How do I solve? and Which method would I use for this one substitution or elimination? I automatically think elimination because of the = sign. Showing work would be helpful, I am printing and using as a study guide. This is very frustrating :/ Thank you so much!
5x + 3y = 4
x - 4y = 3

Answer by Cromlix(4381)

5x + 3y = 4 ....(1)
x - 4y = 3 ....(2)
You now multiply line (1) by 4 and line (2) by 3
This is done to remove one of the unknowns namely 'y'
20x + 12y = 16
3x - 12y = 9
Add
23x = 25
x = 25/23
Substitute x = 25/23 into x - 4y = 3 ...(2)
25/23 - 4y = 3
- 4y = 3 - 25/23
Change 3 to 23rds
- 4y = 69/23 -25/23
-4y = 44/23
-y = 44/92
y = - 44/92 = -22/46 = -11/23
x = 25/23 : y = -11/23